OpenStudy (anonymous):

Does anyone know how to solve this absolute value equation? 14-2|-4-2*1.5|

OpenStudy (amistre64):

if it was written clearer maybe...

OpenStudy (amistre64):

-10 is what my guess is

OpenStudy (anonymous):

The answer is 4 but I have no idea on how to solve it

OpenStudy (anonymous):

I got 0

OpenStudy (amistre64):

8*1.5= 12*-2=-24+14=10

OpenStudy (amistre64):

-10 lol...

OpenStudy (anonymous):

ooops that is not a subtract before the absolute value sign its inside but maybe you are supposed to take 14-2= (12) times the absval -4-2= -6*1.5 = 9 nope think there is a symbol missing or something in equation

OpenStudy (amistre64):

-18+14 = -4; thats closer :)